OpenStudy (anonymous):

A tennis player hits a 58.0 {\rm g} tennis ball so that it goes straight up and reaches a maximum height of 7.35 {\rm m} .How much work does gravity do on the ball on the way up?

OpenStudy (jfraser):

work is force*distance.\[W = F*d\] Force is also equal to m*a \[F = m*a\] so work is also equal to \[W = m*a*d\] the mass of the ball must be in kg. The acceleration of the ball is the acceleration due to gravity (9.8m/s^2), and the distance is the height that the ball reaches. Plug and solve